The mean of 7 positive integers is 405. Two of the integers are 103 and 199 and
the other numbers are greater than 199. If all 7 integers are different, what is
the greatest possible value for any of the 7 integers?

Answer by HasanSahin(52)

The mean is the division of the sum of the numbers by 7 as there are 7 numbers.
So let's find the sum of the 7 numbers >> 405*7 = 2835
Let's find the unknown sum of the 5 numbers by subtracting the sum of the given numbers from the total sum >> 2835-(103+199)=2533
The smallest number of this 5 numbers set must be greater than 199.If we want to find the possible greatest number in this 5 numbers set,we must choose the other 4 numbers minimum.
So I'll choose such that, as they must be different from eachother ;
200,201,202,203 and sum up : 200+201+202+203 = 806
The possible greatest one is 2533-806 = 1727
RF.